OpenStudy (anonymous):

find the unit price for each option shown below round to the nearest cent when necessary indicate which option is the better buy option I: 10 candy bars for 6.75 Option II: 12 candy bars for 7.25 show your work plz or just say the answer

OpenStudy (jack1):

if 10 candy bars cost 6.75 then 1 candy bar costs 6.75/10 = ...?? @I.onlyaskformath.help ?

OpenStudy (anonymous):

10 candy bars for 6$ 12 candy bars for 7$

OpenStudy (anonymous):

What is not to understand? @Jack1

OpenStudy (jack1):

not what i asked, i wanted u to complete the equation 6.75/10 = ...?? =0.675

OpenStudy (jack1):

2nd half if 12 candy bars cost 7.25 then 1 candy bar costs 7.25/12 = ...?? @I.onlyaskformath.help ?

OpenStudy (anonymous):

7.25/12= 0.6041666666666667 @Jack1

OpenStudy (jack1):

yep, so which option is cheaper per bar?

OpenStudy (anonymous):

the 7.25 right?

OpenStudy (jack1):

yep, about 7 cents per bar cheaper

OpenStudy (anonymous):

Thank you

OpenStudy (jack1):

all good dude